# The Flint Hills Series, Mixed Tate Motives, and a Criterion for the Irrationality Measure of $\pi$

### 10 Mar 2026 v1

@@ 6.2. Convergence of the companion series
The rapid decay of increments provides strong numerical evidence for convergence. The convergence is consistent with an effective irrationality measure $\mu(\pi)<5/2$ for the integers sampled.
@@ 6.4. Numerical value of $R_{1}^{*}$ and $S$
$R_{1}^{*}\;\approx\;86.13534,\qquad S\;\approx\;\tfrac{4}{3}\,\zeta(3)+\tfrac{1}{3}\times 86.13534\;\approx\;30.31452.$
@@
Note that $S\approx 30.3$ is the value of the partial series up to $N=10^{5}$, not a quantity near $78$. The value $78.11\ldots$ appearing in the original version of this work corresponded to the intermediate regulator $R_{1}:=\frac{\pi}{2\sqrt{3}}R_{1}^{*}\approx\frac{\pi}{2\sqrt{3}}\times 86.135\approx 78.116$, which is a rescaled version of $R_{1}^{*}$. The two formulations are entirely equivalent; the present paper uses $R_{1}^{*}$ directly to avoid introducing unnecessary scaling factors.
@@ 7.1. The status of $\mu(\pi)\leq 5/2$
Theorem 1.3 reduces the convergence of the Flint Hills series to the Diophantine question $\mu(\pi)\leq 5/2$. The best unconditional bound $\mu(\pi)\leq 7.103$ [16] leaves a substantial gap. Methods that have improved bounds on $\mu(\pi)$ include the Padé approximation approach of Beukers [5] and the integral method of Hata [13]; improvements to these requiring only $\mu(\pi)<5/2$ would resolve the convergence of $S$.
@@ 7.2. Future directions
    Explicit motivic generators. A precise construction of the Mixed Tate Motive $\mathscr{M}$ for which $R_{1}^{*}$ is a period requires specifying an algebraic cycle in $K_{5}(\mathcal{O}_{K})$. This would make Theorem 1.4 fully unconditional within the Beilinson–Bloch framework.
@@
    Effective bounds from motivic structure. If the motivic identity of Theorem 1.4 can be made effective (producing explicit constants $C_{1},C_{2}$), it may be possible to use it to derive new lower bounds on $|n-k\pi|$, contributing to an improvement of $\mu(\pi)$.

### 11 Mar 2026 v2

@@ 6.2. Convergence of the companion series
The rapid decay of increments provides strong numerical evidence for convergence, consistent with an effective irrationality measure $\mu(\pi)<5/2$ for the integers sampled.
@@ 6.3. Transcendental constants
The value of $L(3,\chi_{-3})$ was independently verified via the closed form $4\pi^{3}/(81\sqrt{3})$, via the polygamma formula $(\psi_{2}(2/3)-\psi_{2}(1/3))/54$, and by direct summation of the Dirichlet series; all three agree to 50 decimal places.
@@ 6.4. Numerical value of $R_{1}^{*}$ and $S$
$R_{1}^{*}\approx 86.13534,\qquad S\approx\tfrac{4}{3}\zeta(3)+\tfrac{1}{3}\times 86.13534\approx 30.31451.$
@@
In an earlier version of this work an intermediate quantity $R_{1}:=\pi^{2}\sqrt{3}/(R_{1}^{*})\cdot c$ appeared with numerical value near $78.116$. That quantity arose from a different normalisation of the companion series and is not used in the present formulation. The two formulations are related by a multiplicative rescaling factor of approximately $0.9069$; the present paper works directly with $R_{1}^{*}$ to avoid introducing unnecessary scaling.
@@ 7.1. The status of $\mu(\pi)\leq 5/2$
Theorem 1.3 reduces the convergence of the Flint Hills series to the Diophantine question $\mu(\pi)\leq 5/2$. The best unconditional bound, due to Zeilberger–Zudilin [18], is $\mu(\pi)\leq 7.103205334137$, leaving a substantial gap. Methods that have improved bounds on $\mu(\pi)$ include the Padé approximation approach of Beukers [5] and the integral method of Hata [13]; improvements requiring only $\mu(\pi)<5/2$ would resolve the convergence of $S$.
@@ 7.2. Future directions
    Explicit motivic generators. A precise construction of the Mixed Tate Motive $M$ for which $R_{1}^{*}$ is a period requires specifying an algebraic cycle in $K_{5}(\mathcal{O}_{K})$. This would give the explicit constants $c_{1},c_{2}$ of Conjecture 5.3 and make Theorem 1.4 fully unconditional within the Beilinson–Bloch framework.
@@
    Effective bounds from motivic structure. If the motivic identity of Conjecture 5.3 can be made effective, it may yield new lower bounds on $|n-k\pi|$, contributing to an improvement of $\mu(\pi)$.